Use the method of variation of parameters to solve the system

𝑋′ =[

3 −1 −1

−2 3 2

4 −1 −2

]𝑋+[

1

𝑒𝑡

𝑒𝑡]


1
Expert's answer
2021-09-21T11:46:21-0400

The given question is not at all clear. So, I am taking a example and solving that.


Question: Find a particular solution of:

"\mathbf{y}^{\prime}=\left[\begin{array}{ccc}
2 & -1 & -1 \\
1 & 0 & -1 \\
1 & -1 & 0
\end{array}\right] \mathbf{y}+\left[\begin{array}{c}
e^{t} \\
0 \\
e^{-t}
\end{array}\right]"

Solution:

The characteristic polynomial of the coefficient matrix is

"\left|\begin{array}{lll}
2-\lambda & -1 & -1 \\
1 & -\lambda & -1 \\
1 & -1 & -\lambda
\end{array}\right|=-\lambda(\lambda-1)^{2}"


"\mathbf{y}_{1}=\left[\begin{array}{l}
1 \\
1 \\
1
\end{array}\right],\quad \mathbf{y}_{2}=\left[\begin{array}{l}
e^{t} \\
e^{t} \\
0
\end{array}\right],\quad \text { and } \quad \mathbf{y}_{3}=\left[\begin{array}{l}
e^{t} \\
0 \\
e^{t}
\end{array}\right]"

are linearly independent solutions.


"Y=\left[\begin{array}{lll}
1 & e^{t} & e^{t} \\
1 & e^{t} & 0 \\
1 & 0 & e^{t}
\end{array}\right]" is a fundamental matrix. We seek a particular solution


"\mathbf{y}_{p}=Y \mathbf{u}" where "Y \mathbf{u}^{\prime}=\mathbf{f}" ; that is:

"\left[\begin{array}{lll}
1 & e^{t} & e^{t} \\
1 & e^{t} & 0 \\
e & 0 & e^{t}
\end{array}\right]\left[\begin{array}{l}
u_{1}^{\prime} \\
u_{2}^{\prime} \\
u_{3}^{\prime}
\end{array}\right]=\left[\begin{array}{l}
e^{t} \\
0 \\
e^{-t}
\end{array}\right]"

The determinant of "Y"  is the Wronskian:

"\left|\begin{array}{lll}
1 & e^{t} & e^{t} \\
1 & e^{t} & 0 \\
1 & 0 & e^{t}
\end{array}\right|=-e^{2 t}"

Thus, by Cramer's rule,



"\begin{aligned}
u_{1}^{\prime} &=-\frac{1}{e^{2 t}}\left|\begin{array}{lll}
e^{t} & e^{t} & e^{t} \\
0 & e^{t} & 0 \\
e^{-t} & 0 & e^{t}
\end{array}\right| &=-\frac{e^{3 t}-e^{t}}{e^{2 t}}=e^{-t}-e^{t} \\
u_{2}^{\prime} &=-\frac{1}{e^{2 t}}\left|\begin{array}{ccc}
1 & e^{t} & e^{t} \\
1 & 0 & 0 \\
1 & e^{-t} & e^{t}
\end{array}\right| &=-\frac{1-e^{2 t}}{e^{2 t}}=1-e^{-2 t}
\end{aligned}"

"u_{3}^{\prime}=-\frac{1}{e^{2 t}}\left|\begin{array}{ccc}
1 & e^{t} & e^{t} \\
1 & e^{t} & 0 \\
1 & 0 & e^{-t}
\end{array}\right|=\frac{e^{2 t}}{e^{2 t}} \quad=1"

Therefore


"\mathbf{u}^{\prime}=\left[\begin{array}{l}
e^{-t}-e^{t} \\
1-e^{-2 t} \\
1
\end{array}\right]"

Integrating and taking the constants of integration to be zero yields

"\mathbf{u}=\left[\begin{array}{l}
-e^{t}-e^{-t} \\
\frac{e^{-2 t}}{2}+t \\
t
\end{array}\right] \\
 

 
\mathbf{y}_{p}=Y \mathbf{u}=\left[\begin{array}{lll}
1 & e^{t} & e^{t} \\
1 & e^{t} & 0 \\
1 & 0 & e^{t}
\end{array}\right]\left[\begin{array}{l}
-e^{t}-e^{-t} \\
\frac{e^{-2 t}}{2}+t \\
t
\end{array}\right]=\left[\begin{array}{c}
e^{t}(2 t-1)-\frac{e^{-t}}{2} \\
e^{t}(t-1)-\frac{e^{-t}}{2} \\
e^{t}(t-1)-e^{-t}
\end{array}\right]" is a particular solution.